OpenStudy (anonymous):

derivative of g(t)=5/(sqrt(t))

OpenStudy (anonymous):

use the power rule

OpenStudy (anonymous):

this is a power function

OpenStudy (anonymous):

\[\frac{ 5 }{ \sqrt{t} }\] right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

yeah the power rules

OpenStudy (anonymous):

or you can change it into this \[5t ^{\frac{ -1 }{ 2 }}\]

OpenStudy (anonymous):

that's much more eazier

OpenStudy (anonymous):

\[\large g(t)=\frac{ 5 }{ t^{\frac{ 1 }{ 2 }} }\] \[\large g(t)=5t^{-{\frac{ 1 }{ 2 }}}\]

OpenStudy (anonymous):

I must have missed class that day out sick test tomorrow oh well